# Investigating Resource-efficient Neutron/Gamma Classification ML Models Targeting eFPGAs

### v1

@@ Abstract
There has been considerable interest and resulting progress in implementing machine learning (ML) models in hardware over the last several years from the particle and nuclear physics communities. A big driver has been the release of the Python package, hls4ml, which has enabled porting models specified and trained using Python ML libraries to register transfer level (RTL) code. So far, the primary end targets have been commercial FPGAs or synthesized custom blocks on ASICs. However, recent developments in open-source embedded FPGA (eFPGA) frameworks now provide an alternate, more flexible pathway for implementing ML models in hardware. These customized eFPGA fabrics can be integrated as part of an overall chip design. In general, the decision between a fully custom, eFPGA, or commercial FPGA ML implementation will depend on the details of the end-use application. In this work, we explored the parameter space for eFPGA implementations of fully-connected neural network (fcNN) and boosted decision tree (BDT) models using the task of neutron/gamma classification with a specific focus on resource efficiency. We used data collected using an AmBe sealed source incident on Stilbene, which was optically coupled to an OnSemi J-series SiPM to generate training and test data for this study. We investigated relevant input features and the effects of bit-resolution and sampling rate as well as trade-offs in hyperparameters for both ML architectures while tracking total resource usage. The performance metric used to track model performance was the calculated neutron efficiency at a gamma leakage of 10-3. The results of the study will be used to aid the specification of an eFPGA fabric, which will be integrated as part of a test chip.
@@ 1 Introduction
For a variety of radiation detection applications including neutron radiography [1] [2], associated particle imaging (API) [3] and neutron scatter cameras [4] [5] [6], there is a demonstrated need for compact front-end electronics that are easily scalable to perform individual read out of hundreds to thousands of silicon photomultipliers (SiPMs). Such electronics need to minimize power consumption in order to enable portable deployment. In addition, a requirement for the electronics in the majority of these target applications is an efficient method for discriminating neutron energy depositions from gamma depositions. Ideally, such functionality is implemented at the channel level. The added complexity for this task is that gamma dominated operating environments necessitate very low false positive rates while still maintaining high neutron efficiency.
@@
This task of neutron/gamma classification is fundamentally multi-dimensional. Generally, specific plastic or organic crystalline scintillators with demonstrated inherent pulse shape discrimination (PSD) capability are used. This capability derives from the differences of a scintillator’s singlet/triplet state activation with regards to the type of incident ionizing radiation [7]. However, other factors such as output scintillation wavelength and expected light yield also affect the overall discrimination power [8]. In addition, the time distribution of scintillation light produced is convolved with the SiPM single photon response to produce an output electronic current pulse, introducing yet another dimension. The classification task itself is typically performed by directly or indirectly using differences in the waveform tail between neutron and gamma energy depositions. The standard discriminator used is a ratio between partial and total integration of the waveform. However, several other methods have been utilized, mapping the discrimination problem into different parameter spaces covering both time and frequency domains. Examples include but are not limited to [9] [10] [11] [12]. The majority of such methods do not have straightforward real-time compatible implementations.
@@
For this work, we use Stilbene [13], a widely utilized organic scintillator, and the OnSemi J-series 60035 SiPM [14] to benchmark performance. While the SiPM single photon response is primarily set by details of the chosen SiPM, it is additionally shaped by any series resistance in the signal path and by the bandwidth of front-end electronics used [15]. Although the details of specific front-end topologies are outside the scope of this work, we nevertheless indirectly explore the effects of front-end bandwidth in the study.
Application specific integrated circuits (ASICs) provide a straightforward path to integrating individual read out of large numbers of channels, each with PSD capability, while minimizing power consumption. We have previously investigated one method of performing PSD using only analog circuits when targeting ASICs [16]. Briefly, the method utilizes the mapping of a traditional ratio of partial over total integrated charge to the difference between an event’s scaled total integration value and corresponding partial integration value. A significant benefit results in that PSD capability increases at larger energy depositions rather than remaining flat as a function of energy past a certain energy deposition threshold. This is explicitly shown in Figure 8 of [16]. For an ASIC with a large number of integrated channels, the straight-forward physical circuit implementation requires multiple stages of circuitry at the channel level. In total, this would result in significant area utilization without spending considerable effort on details of circuit implementation. More realistically, a significant area reduction would likely require a mixed-signal solution with the subtraction implemented digitally. However, if already moving towards a mixed-signal solution, we can go beyond a simple translation of the required subtraction operation to a synthesized digital block on-chip. A novel strategy can be used to shift the classification task to a pipelined machine learning (ML) model implemented on an custom embedded field programmable gate array (eFGPA) fabric, which is then integrated on-chip.
@@
This strategy has become viable in the last few years and is explained in more detail in Section 2. In Section 3, we provide a detailed overview of the study conducted in this work to explore the feasibility of an actual implementation, including setup and methodology. We introduce the test bed used for collecting data, provide an overview of the relevant hyperparameters for both types of ML architectures investigated in this study, and detail the metric used for evaluating model performance. In Section 4, we look at the results of the study individually for both ML model architectures. Finally, in Section 5, we summarize the results and its significance for an eventual eFPGA fabric implementation.
@@ 2 The Python to RTL Workflow
The implementation of neutron/gamma classification ML models on commercial FPGAs is well-established. Previous examples of work done include but are not limited to [17] [18] [19] [20] [21] [22]. Shifting from commercial FPGAs to eFPGA fabric targets is enabled by recent developments in both open-source embedded FPGA (eFPGA) frameworks [23] [24] and in porting machine learning (ML) models trained using Python ML libraries to register transfer level (RTL) code [25] [26]. The combined advances in both areas now provide the basis for a viable pathway to a real-time, low-power and compact neutron/gamma classification method using ML models directly on a custom ASIC that integrates all required circuitry. Depending on the exact input features used, analog circuits or digital signal processing algorithms can be used to extract input feature values for a given event. The diagrams in Figure 1 present a generic system signal progression diagram comparing both scenarios. The exact input features required and a decision on if the ability to redefine the input features used is desired will guide such a decision. A re-programmable eFPGA fabric will need to be sufficiently specified to enable such an ability. In either case, the list of circuitry includes necessary front-end electronics for direct SiPM read out, ADCs for either waveform or individual input feature digitization and TDCs for capturing timing information in addition to dedicated digital blocks and/or an appropriately specified eFPGA fabric. A more detailed look at differences between using a waveform digitizer or analog circuits is given in Section 3 using a comparative look at the performance of the standalone charge comparison PSD method. Regardless, it is clear that the specifications of the ADC in both scenarios will directly set the best possible input feature value quantization. In other words, the ADC specification can play a significant role on achievable model performance. To account for this, we chose to fold in the ADC bit-resolution and sampling rate as parameters in our analysis to understand ADC requirements for a future test chip.
In this work, we focus on the first step towards a final integration of a customized open-source eFPGA fabric into an overall chip design for the task of neutron/gamma classification: establishing feasibility. Utilizing an open-source eFPGA framework allows for an FPGA fabric to be integrated as part of an overall chip design without relying on expensive commercial eFPGA vendors. These frameworks allow for customization of the quantity and placement (location) of the various primitive resources (also known as tiles) that make up the fabric, chosen from those available in the open-source framework [23]. This allows for area-efficient solutions with the eFPGA fabric sized for a specific target application. Unlike a commercial FPGA, the availability of specific resource blocks, including memory resources, and their capabilities will depend on both the framework used and the technology node selected.
Achievable area-optimization for the eFPGA fabric will be limited by achievable routing densities and resource capabilities of the typically older process nodes used for the mixed-signal integrated chips targeting the applications under consideration. Explicitly, an eFPGA fabric implemented in 28nm versus 180nm will have very different capabilities for the same arrangement of primitive resource tiles defined using register transfer level (RTL) code. As a result, the exact latencies, resource usage, required maximum clock frequency, etc will be different for the same ML model depending on the process node selected for implementation. Due to these considerations, we used an older AMD/Xilinx Artix 7-series device (xc7a35ticsg324-1L) on an Arty development board as an initial stand-in for this exploration study. This commercial FPGA target has a total of 90 DSP blocks, 41.6k flip-flops, 20.8k LUTs and the equivalent of 100 18Kb RAM blocks. Compared to the top of the line for the 7-series FPGAs, this is an approximately 20x reduction in resources available. Compared against the top of the line UltraScale + family devices, this represents an approximately 60x reduction in available flip-flops and LUTs, 100x reduction for DSPs and 35x reduction for memory resources. The 7-series family devices are fabricated in a 28nm process node, representing a realistic upper bound comparison point on achievable performance and resource capabilities for the expected highest end node targeted for an actual custom eFPGA fabric.
The exact resource usage numbers and other specifications achieved for the stand-in commercial FPGA target will not directly translate for an eFPGA fabric target due to the open source frameworks considered. However, it is expected that the general trends for the models will hold and exact numbers will be off by some scaling factor. We aim to use the results of this study to inform the specification of resources needed for an actual eFPGA fabric to be implemented on a test chip.
In this work, we explored the design parameter space for two basic ML algorithms: boosted decision tree (BDT) and fully connected neural network (fcNN) models. The costs associated with fabricating chips scale proportionally with area for multi-project wafer (MPW) runs. As a result, we specifically investigated feasibility for resource-efficient implementations of these basic ML models, which will translate to an area-efficient implementation of an appropriately sized custom eFPGA fabric to run these models on a test chip.
@@
The traditional FPGA implementation workflow uses RTL code, which specifies the intended behavioral functionality of a digital block, for design entry. From there, (vendor-specific) logic synthesis and place and route (PnR) software tools map the RTL code to resources such as look-up tables (LUTs), flip-flops (FFs), block random access memory (BRAM), etc on a commercial FPGA target die. In the last two decades, commercial high level synthesis (HLS) software has matured and been adopted to allow for bypassing the requirement to write RTL code directly. These tools use C/C++/SystemC/MATLAB code for design entry. The initial code provides an untimed, functional description of the intended digital block behavior and is then automatically mapped to a data flow graph (DFG), from which operations are scheduled based on capabilities of the target process node. The final output of the HLS software is (timed) RTL code.
In order to generally implement ML algorithms in hardware, ML models defined using Python ML libraries need to be mapped to synthesizable C/C++/SystemC/MATLAB code. This last bridge was crossed in the past decade with the recent development of the hls4ml Python package. A detailed explanation of the hls4ml workflow is given in [25]. This package can directly interface with commercial HLS tools such as AMD/Xilinx Vivado HLS and Intel/Altera Quartus HLS to perform high level synthesis on its output C++ code, which functionally describes the ML model.
@@
Originally, support for BDTs was integrated directly into hls4ml along with the support for NN-based ML architectures [26]. Since then, the BDT workflow has been moved into conifer, a complementary Python package released by the same collaboration. The packages, hls4ml and conifer respectively allow specification of quantization levels for the biases/weights or thresholds of the fcNN/BDT model and provides support for pipelining operations. Because of this, the ML model can accept new inputs after waiting a set number of clock cycles (initiation interval), potentially before the final prediction for the previous inference is available (achievable latency). This capability is crucial for replacing PSD circuitry of a large number of individual channels with a single pipelined ML model implemented on an eFPGA fabric. The hls4ml package is flexible enough that it provides several user-configurable settings, which can be set as needed to explore trade-offs in latency, throughput, power and resource usage for the classification task at hand. In comparison, conifer does not provide as extensive configurability, but the design parameter space for BDTs is also simpler [26].
Implementing ML models in hardware requires consideration of several new factors that are generally not necessary for a software implementation. The quantization of the input features plays a significant role in lowering overall resource usage as keeping native 32-bit resolution leads to larger resource usage for not much better performance. For our end-use target application(s), as mentioned, we anticipate having on-chip ADCs. The specifications of the ADC’s bit resolution and full scale range will set the upper bound on the input feature quantization. The process of assigning ADC counts a physical value allows for selecting the base unit, which plays a non-trivial role in the required number of bits per input feature. For example, using millivolts (mV) or Volts (V) for storing the value of a pulse’s peak amplitude is straightforward and interchangeable in software. However, both hls4ml and conifer use fixed point format to represent values for input feature. Thus, the number of integer and fractional bits explicitly need to be set. An additional bit will be used for the sign. Figure 2 explicitly demonstrates the difference in input word lengths required to exactly represent a value over a 2V full-scale range depending on if mV or V is the base unit. We expand on this point and the assumptions used in this study in Section 3.
@@
To explore the full parameter space, we had to pick specific Python libraries to specify and train the ML models. While both scikit-learn and XGBoost have Python libraries allowing for the specification and training of BDTs, it is generally accepted that XGBoost’s version results in both better performance and faster training. Furthermore, preliminary checks of resource usage for similarly set hyperparameter values between BDT models specified using the two libraries indicate very similar resource usage values. For these reasons, we used XGBoost’s BDT model for this work. For fcNNs, several Python libraries are also available. However, at present, only Keras supports a complementary Python library for performing quantization aware training (QKeras), which is also supported by hls4ml. Taking into consideration the importance of this step in the overall aim of resource-efficient ML eFPGA implementations, we have used (Q)Keras for specifying the fcNN models.
@@
The overall requirements on neutron/gamma classification must also take into account the expected range of operating environments for the end-use applications. More specifically, the expected range of both gamma and neutron rates encountered will directly affect the absolute efficiency requirement on the PSD circuitry. We can tolerate larger gamma leakage values if the gamma rates will be less than or equal to neutron rates in the range of expected operating environments. However, in gamma dominated operating environments, much lower gamma leakage values are required. In this work, we assume operation in a gamma-dominated environment and define the performance metric used to assess performance of trained machine learning (ML) models as the neutron efficiency at a gamma leakage of 10-3. This is not an absolute lower bound on expected gamma leakage requirements across all potential end-use applications. Rather, this is due to a statistical limitation from the size of the training data set used for the study as explained further in Section 3.

### v2

@@ Abstract
There has been considerable interest and resulting progress in implementing machine learning (ML) models in hardware over the last several years from the particle and nuclear physics communities. A big driver has been the release of the Python package, hls4ml, which has enabled porting models specified and trained using Python ML libraries to register transfer level (RTL) code. So far, the primary end targets have been commercial field-programmable gate arrays (FPGAs) or synthesized custom blocks on application specific integrated circuits (ASICs). However, recent developments in open-source embedded FPGA (eFPGA) frameworks now provide an alternate, more flexible pathway for implementing ML models in hardware. These customized eFPGA fabrics can be integrated as part of an overall chip design. In general, the decision between a fully custom, eFPGA, or commercial FPGA ML implementation will depend on the details of the end-use application. In this work, we explored the parameter space for eFPGA implementations of fully-connected neural network (fcNN) and boosted decision tree (BDT) models using the task of neutron/gamma classification with a specific focus on resource efficiency. We used data collected using an AmBe sealed source incident on Stilbene, which was optically coupled to an OnSemi J-series silicon photomultiplier (SiPM) to generate training and test data for this study. We investigated relevant input features and the effects of bit-resolution and sampling rate as well as trade-offs in hyperparameters for both ML architectures while tracking total resource usage. The performance metric used to track model performance was the calculated neutron efficiency at a gamma leakage of 10-3. The results of the study will be used to aid the specification of an eFPGA fabric, which will be integrated as part of a test chip.
@@ 1 Introduction
The initial release and continued development of the hls4ml Python package over the past several years has enabled a complete Python to hardware pipeline for fast deployment of neural network (NN) based machine learning (ML) models. A detailed explanation of the hls4ml workflow is given in [1] and [2]. This package provides front-end converters to map models defined using standard Python ML libraries to an internal hls4ml representation and then, after certain optimizations, directly outputs synthesizable C++ code that functionally describes the ML model. The package targets the input requirements for specific commercial high level synthesis (HLS) tools such as AMD/Xilinx Vivado HLS [3] and Intel/Altera Quartus HLS [4] by using the appropriate C++ HLS libraries and pragmas/directives. Originally, support for boosted decision trees (BDTs) was integrated directly into hls4ml along with the support for NN-based ML architectures [5]. Since then, the BDT workflow has been moved into conifer, a complementary Python package released by the same collaboration. The original hardware target for these packages were commercial field-programmable gate array (FPGA) devices. However, over the years, the packages have been successfully used for fully custom ML model implementation on application specific integrated circuits (ASICs) [6] and even more recently, embedded FPGA (eFPGA) fabrics designed using open-source frameworks [7].
This latest hardware target was made possible by recent developments of open-source eFPGA frameworks, such as FABulous [8] and OpenFPGA [9], which has removed the need to rely on expensive commercial vendors to integrate FPGA fabrics as part of an overall chip design. We intend to use FABulous, which allows for customization of the quantity and placement (location) of various primitive resources (also known as tiles) that will make up the fabric, chosen from those available in the framework. In addition, it provides the flexibility to implement custom tiles that can be integrated as part of an overall fabric. Overall, this technology allows for area-efficient solutions by specifying and sizing the eFPGA fabric for a given target application. However, there are several new considerations to using this technology. A detailed overview of the fabric generation workflow and these considerations are given in [7] and are summarized and expanded on here. Unlike a commercially available FPGA device, the availability of specific resource blocks, including memory resources, and their capabilities will depend on the framework and the technology node selected for physical implementation of the designed fabric. In addition, these resources must be explicitly included in the custom eFPGA fabric for them to be available.
[…]
A hardware-aware co-design methodology can aid in sparsifying ML models to target eFPGA deployments. For particle and nuclear physics applications, the eFPGA fabric is expected to be a part of an overall chip design, which will include dedicated front-end circuits to read out sensors, data converters, and digital back-ends to interface with data acquisition boards. By integrating data converter specifications into the ML model specification and training, we can trade-off the area/resource requirements of the eFPGA fabric with final converter specifications to deploy a specific ML model while maintaining the necessary level of performance. In addition, deployment of ML models on custom eFPGA fabrics means that we can also update the models as needed over the lifetime of the experiment to account for changes in detector conditions, different operations modes (i.e. calibration runs), etc on top of the benefit to ML-aided solutions to optimizing data to save for further offline analysis.
Because the number of resources (physical area) required to deploy ML models on eFPGAs will directly determine feasibility, this work explored the design parameter space (targeting eFPGA deployments) for two very basic ML algorithms: boosted decision tree (BDT) and fully connected neural network (fcNN) models. More specifically, we investigated feasibility for very sparse implementations of these basic ML models, which will translate to an area-efficient implementation for an appropriately sized custom eFPGA fabric to run these models on a future test chip.
We used an older AMD/Xilinx Artix 7-series [11] device (xc7a35ticsg324-1L) on an Arty development board as an initial stand-in for this exploration study. The exact resource usage numbers and other specifications achieved with this study using the stand-in commercial FPGA target will not directly translate to an open-source framework designed custom eFPGA fabric target. However, it is expected that the general trends for the models will hold and exact numbers will be off by some scaling factor. The commercial FPGA target has a total of 90 digital signal processing (DSP) blocks, 41.6k flip-flops (FFs), 20.8k lookup tables (LUTs) and the equivalent of 100 random-access memory (RAM) blocks. Compared to the top of the line for the 7-series FPGAs, this is an approximately 20 reduction in resources available. Compared against the top of the line UltraScale+ family devices [12], this represents an approximately 60 reduction in available flip-flops and LUTs, 100 reduction for DSPs and 35 reduction for memory resources. The 7-series family devices are fabricated in a process node, representing a realistic upper bound comparison point on achievable performance and resource capabilities for the expected highest end node targeted for an actual custom eFPGA fabric. The amount of resources available on this commercial device is expected to be higher than for a reasonable custom FABulous eFPGA fabric, especially for the older process nodes. Thus, a main focus of this work was to target using only a small fraction of the overall resources available on this device when specifying a high performance ML model. We shall expand on the details in the next section.
@@
To frame this feasibility study, the non-trivial task of neutron/gamma classification was used as an initial case study. We aim to use the results of this study to inform the specification of an actual eFPGA fabric to be implemented on the test chip, which will be able to deploy high performance ML models for this classification task. The task of neutron/gamma classification is either explicitly required or desired in a variety of radiation detection applications including neutron radiography [13] [14], associated particle imaging (API) [15] and neutron scatter cameras [16] [17] [18]. These applications typically happen in gamma dominated operating environments, necessitating very low false positive rates while still maintaining high neutron efficiency. The task is fundamentally multi-dimensional. Generally, specific plastic or organic crystalline scintillators with demonstrated inherent pulse shape discrimination (PSD) capability are used. This capability derives from the differences of a scintillator’s singlet/triplet state activation with regards to the type of incident ionizing radiation [19]. However, other factors such as output scintillation wavelength and expected light yield also affect the overall discrimination power [20]. In addition, the time distribution of scintillation light produced is convolved with the silicon photomultiplier (SiPM) single photon response to produce an output electronic current pulse, introducing yet another dimension. The classification task itself is typically performed by directly or indirectly using differences in the waveform tail between neutron and gamma energy depositions. The standard discriminator used is a ratio between partial and total integration of the waveform.
The implementation of neutron/gamma classification ML models on commercial FPGAs is established. Previous examples of work done include but are not limited to [21] [22] [23] [24] [25] [26]. However, the combined advances in both open-source eFPGAs and translation of ML models to hardware now provide the basis for a viable pathway to a real-time, low-power and compact neutron/gamma classification method using ML models directly on a custom ASIC that integrates all required circuitry, including the eFPGA fabric. Thus, it is important to note that while this classification task is used as a case study for this work, a successful future eFPGA demonstration would open up new possibilities in implementing the required discrimination capability for the target applications.
@@
For this work, we use Stilbene [27], a widely utilized organic scintillator, and the OnSemi J-series 60035 SiPM [28] to benchmark performance. While the SiPM single photon response is primarily set by details of the chosen SiPM, it is additionally shaped by any series resistance in the signal path and by the bandwidth of front-end electronics used [29]. Although the details of specific front-end topologies are outside the scope of this work, we nevertheless indirectly explore the effects of front-end bandwidth in the study through data converter specifications.
@@
Further details on using hls4ml and conifer to deploy ML models in hardware and the details of the workflow are given in Section 2. In Section 3, we provide a detailed overview of the study conducted in this work to explore the feasibility of a high performance eFPGA targeted ML model implementation, including setup and methodology. We introduce the test bed used for collecting data, provide an overview of the relevant hyperparameters for both types of ML architectures investigated in this study, and detail the metric used for evaluating model performance. In Section 4, we look at the results of the study individually for both ML model architectures. Finally, in Section 5, we summarize these results and its significance for an eventual embedded field-programmable gate array (eFPGA) fabric implementation.
@@ 2 The Python to RTL Workflow
The traditional FPGA implementation workflow uses RTL code, which specifies the intended behavioral functionality of a digital block, for design entry. From there, (vendor-specific) logic synthesis and place and route (PnR) software tools map the RTL code to resources such as LUTs, FFs, block RAM (BRAM), etc on a commercial FPGA target die. In the last two decades, commercial HLS software has matured and been adopted to allow for bypassing the requirement to write RTL code directly [30]. These tools use C/C++/SystemC/MATLAB code for design entry. The initial code provides an untimed, functional description of the intended digital block behavior and is then automatically mapped to a data flow graph (DFG), from which operations are scheduled based on capabilities of the target process node. The final output of the HLS software is (timed) RTL code that can be used in the standard FPGA workflow.
@@
As mentioned, the packages, hls4ml and conifer provide C++ input code for the HLS tools. In addition, they, respectively, allow specification of quantization levels for the biases/weights or thresholds of the fcNN/BDT model and provide support for pipelining operations. Because of this, the ML model can accept new inputs after waiting a set number of clock cycles (initiation interval), potentially before the final prediction for the previous inference is available (achievable latency). This capability is crucial for potentially replacing neutron/gamma classification circuitry of a large number of individual channels with a single pipelined ML model implemented on an eFPGA fabric. The hls4ml package is flexible enough that it provides several user-configurable settings, which can be set as needed to explore trade-offs in latency, throughput, power and resource usage for the classification task at hand [2]. In comparison, conifer does not provide as extensive configurability, but the design parameter space for BDTs is also simpler [5].
@@
To explore the full parameter space, we had to pick specific Python libraries to specify and train the ML models. While both scikit-learn [31] and XGBoost [32] have Python libraries allowing for the specification and training of BDTs, it is generally accepted that XGBoost’s version results in both better performance and faster training [32]. Furthermore, preliminary checks of resource usage for similarly set hyperparameter values between BDT models specified using the two libraries indicate very similar resource usage values. For these reasons, we used XGBoost’s BDT model for this work. For fcNNs, several Python libraries are also available. Currently, both Pytorch [33] and TensorFlow [34], which has integrated Keras [35], support performing quantization aware training (QAT). However, hls4ml primarily supports QAT, at present, through the QKeras front-end [36]. Taking into consideration the importance of this step in the overall aim of resource-efficient ML eFPGA implementations [1], we have used QKeras for specifying the fcNN models.
Implementing ML models in hardware requires consideration of several new factors that are generally not necessary for a software implementation. The choice of input features to use in the ML model will depend on the ability to extract these features in real-time. Depending on the exact input features used, analog circuits or digital signal processing algorithms can be used to extract input feature values for a given event. The diagrams in Figure 1 present a generic system signal progression diagram comparing both scenarios. The exact input features required and a decision on if the ability to redefine the input features used is desired will guide such a decision. A re-programmable eFPGA fabric will need to be sufficiently specified to enable such an ability. In either case, the list of circuitry includes necessary front-end electronics for direct SiPM read out, analog-to-digital converters (ADCs) for either waveform or individual input feature digitization and time-to-digital converters (TDCs) for capturing timing information in addition to dedicated digital blocks and/or an appropriately specified eFPGA fabric. A more detailed look at differences between using a waveform digitizer or analog circuits is given in Section 3 using a comparative look at the performance of the standalone charge comparison pulse shape discrimination method. Regardless, it is clear that the specifications of the ADC in both scenarios will directly set the best possible input feature value quantization. In other words, the ADC specification can play a significant role on achievable model performance. To account for this, we chose to fold in the ADC bit-resolution and sampling rate as parameters in our analysis to understand possible trade-offs of ADC requirements for a future test chip (i.e. perform a hardware aware co-design of the fcNN/BDT models).
[…]
@@
The overall requirements on neutron/gamma classification must also take into account the expected range of operating environments for the end-use applications. More specifically, the expected range of both gamma and neutron rates encountered will directly affect the absolute efficiency requirement on the neutron/gamma classification circuitry. We can tolerate larger gamma leakage values if the gamma rates will be less than or equal to neutron rates in the range of expected operating environments. However, in gamma dominated operating environments, much lower gamma leakage values are required. In this work, we assume operation in a gamma-dominated environment and define the performance metric used to assess performance of trained machine learning (ML) models as the neutron efficiency at a gamma leakage of 10-3. This is not an absolute lower bound on expected gamma leakage requirements across all potential end-use applications. Rather, this is due to a statistical limitation from the size of the training data set used for the study as explained further in Section 3.